find the EAR in each of the following cases (use 365 days a year. Do not round intermediate calculations and enter your answers as a percent rounded to 2 decimal places)

started apr 10.0% quarterly
started apr 19% monthly
started apr 15.0%Daily
started apr 12.0%infinite

please show the work not Shure what I'm doing wrong.
Answer #1

1. EAR = (1 + 10%/4)^4 - 1 = 10.38%

2. EAR = (1 + 19%/12)^12 - 1 = 20.75%

3. EAR = (1 + 15%/365)^365 - 1 = 16.18%

4. EAR = e0.12 - 1 = 12.75%